# Copositivity criteria of a class of fourth order 3-dimensional symmetric tensors

### v2

@@ 3.1 Analytical expressions of strict copositivity
    there is at least one $1$ in $\{t_{1122},t_{1133},t_{2233}\}$ if two of $\{t_{1123},t_{1223},t_{1233}\}$ are only $-1$.
Proof. Necessity. If $\mathcal{T}$ is srictly copositive, but the conditions don’t hold. Then for $x=(1,1,1)^{\top}$, it follows that (1) one of $\{t_{1122},t_{1133},t_{2233}\}$ is $-1$,
@@
(2) $t_{1122}=t_{1133}=t_{2233}=-1$,
$\mathcal{T}x^{4}=$ $\sum_{i,j,k,l=1}^{3}t_{ijkl}x_{i}x_{j}x_{k}x_{l}=27-18-24+12=-3<0.$
@@
(2) We might take $t_{1122}=t_{1233}=1$ and $t_{1133}=t_{2233}=t_{1123}=t_{1223}=-1$, then
$\mathcal{T}x^{4}=$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+6x_{1}^{2}x_{2}^{2}-6x_{1}^{2}x_{3}^{2}-6x_{2}^{2}x_{3}^{2}+4x_{1}^{3}x_{2}+4x_{1}^{3}x_{3}+4x_{1}x_{2}^{3}+4x_{1}x_{3}^{3}+4x_{2}^{3}x_{3}+4x_{2}x_{3}^{3}$
$-12x_{1}^{2}x_{2}x_{3}-12x_{1}x_{2}^{2}x_{3}+12x_{1}x_{2}x_{3}^{2},$
@@
    $x_{1}\geq x_{3}$ and $x_{1}>0$, $(\mathcal{T}x^{3})_{1}>0,$ which is done; otherwise,
@@
    two of $\{t_{1123},t_{1223},t_{1233}\}$ are only $-1$ and there is at least one $1$ in $\{t_{1122},t_{1133},t_{2233}\}$;
@@ 3.2 Analytical expressions of copositivity
    there is at least one $1$ in $\{t_{iiij};i,j=1,2,3,i\neq j\}$ if there is only one $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$;
@@
    there is at least two $1$ in $\{t_{iiij};i,j=1,2,3,i\neq j\}$ if there is only two $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$;
@@
$=$ $(3+1-1)^{4}+8(3^{2}-3^{3}-3)<0.$
@@
    there is at least one $1$ in $\{t_{iiij};i,j=1,2,3,i\neq j\}$ if there is only one $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$;
@@
    there is at least two $1$ in $\{t_{iiij};i,j=1,2,3,i\neq j\}$ if there is only two $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$;
@@ 3.3 Applications to a gerenal tensor
It is obvious that the copositivity of symmetric tensor $\mathcal{T}=(t_{ijkl})$ is equivalent to the copositivity of $\mathcal{T^{\prime}}=(t^{\prime}_{ijkl})$. So, applying Corollaries 3.2 and 3.6 (or Theorems 3.4 (1) and 3.8 (1) and (2) to establish easily the following conclusions .
@@
$t_{1123}\geq t_{1111}^{\frac{1}{2}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{4}},t_{1223}\geq t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{2}}t_{3333}^{\frac{1}{4}},t_{1233}\geq t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{2}}$
@@
$t_{iijj}\geq\sqrt{t_{iiii}t_{jjjj}},\ \ t_{iiij}\geq-t_{iiii}^{\frac{3}{4}}t_{jjjj}^{\frac{1}{4}}.$
@@
and there is one $t_{kkll}\in\{t_{iijj};i,j=1,2,3,i\neq j\}$ such that
$t_{kkll}\geq\sqrt{t_{kkkk}t_{llll}}\mbox{ and }t_{iijj}\geq-\sqrt{t_{iiii}t_{jjjj}}\mbox{ for all }t_{iijj}\neq t_{kkll}.$
@@
and there is two $t_{kkkl},t_{sssr}\in\{t_{iiij};i,j=1,2,3,i\neq j\}$ such that
$t_{kkkl}\geq t_{kkkk}^{\frac{3}{4}}t_{llll}^{\frac{1}{4}},\ t_{sssr}\geq t_{sss}^{\frac{3}{4}}t_{rrrr}^{\frac{1}{4}}$
and for all $t_{iiij}\in\{t_{iiij};i,j=1,2,3,i\neq j\}\setminus\{t_{kkkl},t_{sssr}\},$
$t_{iiij}\geq-t_{iiii}^{\frac{3}{4}}t_{jjjj}^{\frac{1}{4}}.$
@@ 3.4 Ternary quartic inequalities
$f(x_{1},x_{2},x_{3})=$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+6x_{1}^{2}x_{2}^{2}+6x_{1}^{2}x_{3}^{2}+6x_{2}^{2}x_{3}^{2}-4x_{1}^{3}x_{2}+4x_{1}^{3}x_{3}-4x_{1}x_{2}^{3}-4x_{1}x_{3}^{3}$
$=$ $(x_{1}-x_{2}+x_{3})^{4}-8x_{1}x_{3}^{2}(x_{3}-3x_{2}).$
From the arguments of Theorems 3.7 (1), it follows that the zero of function $f(x_{1},x_{2},x_{3})$ must be on the boundary of the non-negative orthant (that is, three coordinate planes), and so, $f(x_{1},x_{2},x_{3})=0$ if and only if $x_{1}=x_{2}$, $x_{3}=0$ or $x_{1}=0,$ $x_{2}=x_{3}$.

### v3

@@ 3.1 Analytical expressions of strict copositivity
    there is at least one $1$ in $\{t_{1122},t_{1133},t_{2233}\}$ and $t_{iijk}=t_{ijjk}=-t_{iijj}=-1$ for $i\neq j$ and $k\neq j$ and $i\neq k$ if two of $\{t_{1123},t_{1223},t_{1233}\}$ are only $-1$.
Proof. Necessity. If $\mathcal{T}$ is srictly copositive, but the conditions don’t hold. Then
(1) for $x=(1,1,1)^{\top}$, it follows that one of $\{t_{1122},t_{1133},t_{2233}\}$ is $-1$,
@@
(2) for $x=(1,1,1)^{\top}$, it follows that $t_{1122}=t_{1133}=t_{2233}=-1$,
$=$ $27-18-24+12=-3<0;$
Let $t_{1133}=1,t_{1122}=t_{2233}=-1$, $t_{1233}=1,t_{1223}=t_{1123}=-1$. For $x=(2,2.1,1)^{\top}$,
$\mathcal{T}x^{4}=$ $(x_{1}+x_{2}+x_{3})^{4}-12(x_{1}^{2}x_{2}^{2}+x_{2}^{2}x_{3}^{2})-24(x_{1}^{2}x_{2}x_{3}+x_{1}x_{2}^{2}x_{3})$
$=$ $(2+2.1+1)^{4}-12(2^{2}\times 2.1^{2}+2.1^{2})-24(2^{2}\times 2.1+2\times 2.1^{2})$
$=$ $-1.3599<0.$
@@
(2) We might take $t_{1122}=t_{1133}=t_{1233}=1$ and $t_{2233}=t_{1123}=t_{1223}=-1$, then
$\mathcal{T}x^{4}=$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+6x_{1}^{2}x_{2}^{2}+6x_{1}^{2}x_{3}^{2}-6x_{2}^{2}x_{3}^{2}+4x_{1}^{3}x_{2}+4x_{1}^{3}x_{3}+4x_{1}x_{2}^{3}+4x_{1}x_{3}^{3}+4x_{2}^{3}x_{3}+4x_{2}x_{3}^{3}$
$-12x_{1}^{2}x_{2}x_{3}-12x_{1}x_{2}^{2}x_{3}+12x_{1}x_{2}x_{3}^{2}$
$\geq$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+6x_{1}^{2}x_{2}^{2}-6x_{1}^{2}x_{3}^{2}-6x_{2}^{2}x_{3}^{2}+4x_{1}^{3}x_{2}+4x_{1}^{3}x_{3}+4x_{1}x_{2}^{3}+4x_{1}x_{3}^{3}+4x_{2}^{3}x_{3}+4x_{2}x_{3}^{3}$
$-12x_{1}^{2}x_{2}x_{3}-12x_{1}x_{2}^{2}x_{3}+12x_{1}x_{2}x_{3}^{2}=\mathcal{T}^{\prime}x^{4},$
@@
    $x_{1}\geq x_{3}$ and $x_{1}>0$, $(\mathcal{T}^{\prime}x^{3})_{1}>0,$ which is done; otherwise,
@@
    two of $\{t_{1123},t_{1223},t_{1233}\}$ are only $-1$ and $t_{iijk}=t_{ijjk}=-t_{iijj}=-1$ for $i\neq j$ and $k\neq j$ and $i\neq k$;
@@ 3.2 Analytical expressions of copositivity
Let $\mathcal{T}=(t_{ijkl})$ be a 4th-order $3$-dimensional symmetric tensor with $|t_{ijkl}|=t_{iiii}=t_{iijj}=1$ for all $i,j,k,l\in\{1,2,3\}$. Then $\mathcal{T}$ is copositive if and only if
@@
    there is at least one $1$ in $\{t_{iiij};i,j=1,2,3,i\neq j\}$ and $t_{iijk}=t_{iiij}t_{iiik}=-1$ or $t_{iijk}=-t_{iiij}=-t_{iiik}=-1$for $i\neq j$ and $k\neq j$ and $i\neq k$ if there is only one $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$;
@@
    there is at least two $1$ in $\{t_{iiij};i,j=1,2,3,i\neq j\}$ and there is at least one $1$ in $\{t_{iiij},t_{iiik};t_{iijk}=-1,i\neq j,i\neq k,i\neq k\}$ if there is only two $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$;
@@
$=$ $(3+1-1)^{4}+8(3^{2}-3^{3}-3)<0;$
Assume $t_{iijk}=t_{iiij}=t_{iiik}=-1$. We might take $t_{1113}=t_{1123}=t_{1233}=1$, $t_{1223}=t_{1222}=t_{2223}=t_{1333}=t_{1112}=t_{2333}=-1$, then for $x=(1,3,1)^{\top}$,
$\mathcal{T}x^{4}=$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+6x_{1}^{2}x_{2}^{2}+6x_{1}^{2}x_{3}^{2}+6x_{2}^{2}x_{3}^{2}-4x_{1}^{3}x_{2}+4x_{1}^{3}x_{3}-4x_{1}x_{2}^{3}-4x_{1}x_{3}^{3}$
$-4x_{2}^{3}x_{3}-4x_{2}x_{3}^{3}+12x_{1}^{2}x_{2}x_{3}-12x_{1}x_{2}^{2}x_{3}+12x_{1}x_{2}x_{3}^{2}$
$=$ $(x_{1}+x_{2}-x_{3})^{4}+8(3x_{1}^{2}x_{2}x_{3}-x_{1}^{3}x_{2}-x_{1}x_{2}^{3}+x_{1}^{3}x_{3})$
$=$ $(1+3-1)^{4}+8(3^{2}-3-3^{3}+1)<0.$
@@
Assume $t_{iiij}=t_{iiik}=t_{iijk}=-1.$ We might take $t_{1333}=t_{2333}=t_{1123}=1$ and $t_{1113}=t_{1112}=t_{2223}=t_{1222}=t_{1233}=t_{1223}=-1$. Then for $x=(1,3,1)^{\top}$,
$\mathcal{T}x^{4}=$ $(x_{1}+x_{2}+x_{3})^{4}-8(3x_{1}x_{2}x_{3}^{2}+3x_{1}x_{2}^{2}x_{3}+x_{1}^{3}x_{2}+x_{1}x_{2}^{3}+x_{1}x_{3}^{3}+x_{2}^{3}x_{3})$
$=$ $5^{4}-8(9+27+3+27+1+27)=-127<0;$
@@
    there is only one $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$ and $t_{iijk}=t_{iiij}t_{iiik}=-1$ or $t_{iijk}=-t_{iiij}=-t_{iiik}=-1$for $i\neq j$ and $k\neq j$ and $i\neq k$;
@@
    there is only two $-1$ in $\{t_{1123},t_{1223},t_{1233}\}$ and there is at least one $1$ in $\{t_{iiij},t_{iiik};t_{iijk}=-1,i\neq j,i\neq k,i\neq k\}$;
@@ 3.3 Applications to a gerenal tensor
It is obvious that the copositivity of symmetric tensor $\mathcal{T}=(t_{ijkl})$ is equivalent to the copositivity of $\mathcal{T^{\prime}}=(t^{\prime}_{ijkl})$. So, applying Corollaries 3.2 and 3.6 (or Theorems 3.4 (1) and 3.8 (2) to establish easily the following conclusions .
@@
Let $\mathcal{T}=(t_{ijkl})$ be a 4th-order $3$-dimensional symmetric tensor with $t_{iiii}>0$ for all $i\in\{1,2,3\}$ Assume one of the following three conditions holds,
  • (a)
    $t_{1123}\geq-t_{1111}^{\frac{1}{2}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{4}},t_{1223}\geq t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{2}}t_{3333}^{\frac{1}{4}},t_{1233}\geq t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{2}}$;
  • (b)
    $t_{1123}\geq t_{1111}^{\frac{1}{2}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{4}},t_{1223}\geq-t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{2}}t_{3333}^{\frac{1}{4}},t_{1233}\geq t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{2}}$;
  • (c)
    $t_{1123}\geq t_{1111}^{\frac{1}{2}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{4}},t_{1223}\geq t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{2}}t_{3333}^{\frac{1}{4}},t_{1233}\geq-t_{1111}^{\frac{1}{4}}t_{2222}^{\frac{1}{4}}t_{3333}^{\frac{1}{2}}$.
@@
$t_{iijj}\geq\sqrt{t_{iiii}t_{jjjj}},\ \ t_{iiij}\geq-t_{iiii}^{\frac{3}{4}}t_{jjjj}^{\frac{1}{4}},$
and there is at least one $t_{rrrs}\in\{t_{iiij},t_{iiik};t_{iijk}\geq-t_{iiii}^{\frac{1}{2}}t_{jjjj}^{\frac{1}{4}}t_{kkkk}^{\frac{1}{4}},i\neq j,i\neq k,k\neq j\}$ such that
$t_{rrrs}\geq t_{rrrr}^{\frac{3}{4}}t_{ssss}^{\frac{1}{4}}.$
@@
and
$t_{iijj}\geq\sqrt{t_{iiii}t_{jjjj}}\mbox{ and }t_{iijk}\geq-t_{iiii}^{\frac{1}{2}}t_{jjjj}^{\frac{1}{4}}t_{kkkk}^{\frac{1}{4}}\mbox{ and }t_{ijjk}\geq-t_{iiii}^{\frac{1}{4}}t_{jjjj}^{\frac{1}{2}}t_{kkkk}^{\frac{1}{4}}.$
@@
and there is at least one $t_{sssr}\in\{t_{iiij},t_{iiik};t_{iijk}\geq-t_{iiii}^{\frac{1}{2}}t_{jjjj}^{\frac{1}{4}}t_{kkkk}^{\frac{1}{4}},i\neq j,i\neq k,k\neq j\}$ such that
$t_{sssr}\geq t_{ssss}^{\frac{3}{4}}t_{rrrr}^{\frac{1}{4}}.$
@@ 3.4 Ternary quartic inequalities
$f(x_{1},x_{2},x_{3})=$ $x_{1}^{4}+x_{2}^{4}+x_{3}^{4}+6x_{1}^{2}x_{2}^{2}+6x_{1}^{2}x_{3}^{2}+6x_{2}^{2}x_{3}^{2}-4x_{1}^{3}x_{2}-4x_{1}^{3}x_{3}-4x_{1}x_{2}^{3}-4x_{1}x_{3}^{3}$
$-4x_{2}^{3}x_{3}-4x_{2}x_{3}^{3}+12x_{1}^{2}x_{2}x_{3}+12x_{1}x_{2}^{2}x_{3}+12x_{1}x_{2}x_{3}^{2}$
$=$ $(x_{1}+x_{2}+x_{3})^{4}-8(x_{1}^{3}x_{3}+x_{1}x_{2}^{3}+x_{2}^{3}x_{3}+x_{2}x_{3}^{3}+x_{1}^{3}x_{2}+x_{1}x_{3}^{3}).$
From the arguments of Theorems 3.5, it follows that the zero of function $f(x_{1},x_{2},x_{3})$ must be on the boundary of the non-negative orthant (that is, three coordinate planes), and so, $f(x_{1},x_{2},x_{3})=0$ if and only if $x_{1}=x_{2}$, $x_{3}=0$ or $x_{1}=0,$ $x_{2}=x_{3}$ or $x_{1}=x_{3},x_{2}=0$.